OpenStudy (anonymous):

evaluate the limit: as x-->0, sin (7x)/2x

OpenStudy (nenadmatematika):

7/2

myininaya (myininaya):

\[\frac{\sin(7x)}{7x} \cdot \frac{7}{2}\]

myininaya (myininaya):

that should help let me know if you need more help

OpenStudy (anonymous):

So, is it 7/2 since as x->0 lim sin(x)/(x) =1...I would just do something like lim (7/2) [sin(x)/(x)]?

OpenStudy (nenadmatematika):

yes, you could say so

OpenStudy (anonymous):

Thanks!!